OpenStudy (anonymous):

Determine how long it would take an investment of $10000 to double in value is it were compounded annually at 9.75%

OpenStudy (anonymous):

would it just be 812.5 per month?

hero (hero):

Just use the formula

OpenStudy (anonymous):

what formula? i wasnt given one :S

hero (hero):

Formula for compound interest

OpenStudy (anonymous):

M = P( 1 + i )^n?

hero (hero):

Yeah, but, we usually use A instead of M

OpenStudy (anonymous):

oh ok! Thanks a bunch hero

hero (hero):

Actually, the full formula is \(A = P(1 + \frac{r}{n})^{nt}\) To figure out when it doubles, use: \(2A = P(1 + \frac{r}{n})^{nt}\)

OpenStudy (anonymous):